7(3x + 4) = 8(2x+ 5) + 13

Answer by checkley75(3666)

7(3X+4)=8(2X+5)+13
21X+28=16X+40+13
21X-16X=53-28
5X=25
X=25/5
X=5 ANSWER.
PROOF
7(3*5+4)=8(2*5+5)+13
7(15+4)=8(10+5)+13
7*19=8*15+13
133=120+13
133=133